OpenStudy (anonymous):

What percent of 60 is 12? Show your work.

OpenStudy (anonymous):

20 percent

OpenStudy (anonymous):

x percent . 60 = 12 (x/100).60 =12 60 x = 1200 x = 1200/60 x = 20